# Interpret the estimand framework from a causal inference perspective

### v2

@@ Abstract
Keywords: estimand, causal inference, ICH E9
@@ 1 Introduction
The estimand framework was drafted in 2017 by ICH in Efficacy Guideline E9(R1) as addendum to Efficacy Guideline E9 and later published in 2019 [1, 2]. It has gained increasing attention in the pharmaceutical industry and the academia [4, 5, 3, 6, 7, 8]. Many clinical trials have used the estimand framework to develop new drugs for both oncology and non-oncology diseases, and professional working groups such as the Oncology Estimand Working Group (https://oncoestimand.github.io/oncowg_webpage/docs/) have been initiated to study how the estimand framework should be incorporated in pharmaceutical practices [9].
@@ 3.5 The principal stratum strategy
This strategy estimates the pure treatment effect in a different target population. When specific subpopulations are of clinical interest, this strategy would be appropriate. Statistical methods with regard to this strategy may be complex. Bayesian inference would be a good choice for this strategy [13, 14].
@@ 3.6 Further possibilities
The above five strategies proposed in the estimand framework can be used in a flexible way. Several strategies can be used together. For example, the treatment policy strategy to analyze concomitant therapies can be used with the principal stratum strategy to identify a subpopulation. A composite strategy has been proposed by industrial professionals that the treatment policy strategy to analyze allowed concomitant therapies is combined with the hypothetical strategy to analyze prohibited concomitant therapies [9].
@@ 5 Conclusion
The estimand framework is advantageously useful for estimation of causal treatment effects. Learning and strengthening statistical causal inference theories behind the estimand framework would facilitate implementation and improvement of the estimand framework in industry.

### v3

@@ Abstract
Keywords: Estimand; Causal inference; ICH E9; Clinical trial; Treatment effect; Pharmaceutical; Biomedical research
@@ 1 Introduction
The estimand framework was drafted in 2017 by ICH in Efficacy Guideline E9(R1) as addendum to Efficacy Guideline E9 and later published in 2019 [1, 2]. It has gained increasing attention in the pharmaceutical industry and the academia [3, 4, 5, 6, 7]. Many clinical trials have used the estimand framework to develop new drugs for both oncology and non-oncology diseases, and professional working groups such as the Oncology Estimand Working Group have been initiated to study how the estimand framework should be incorporated in pharmaceutical practices [8].
@@ 3.5 The principal stratum strategy
This strategy estimates the pure treatment effect in a different target population. When specific subpopulations are of clinical interest, this strategy would be appropriate. Statistical methods with regard to this strategy may be complex. Bayesian inference would be a good choice for this strategy [12, 13, 14].
@@ 3.6 Further possibilities
The above five strategies proposed in the estimand framework can be used in a flexible way. Several strategies can be used together. For example, the treatment policy strategy to analyze concomitant therapies can be used with the principal stratum strategy to identify a subpopulation. A composite strategy has been proposed by industrial professionals that the treatment policy strategy to analyze allowed concomitant therapies is combined with the hypothetical strategy to analyze prohibited concomitant therapies.
@@ 5 Conclusion
The estimand framework is advantageously useful for estimation of causal treatment effects. Learning and strengthening statistical causal inference theories behind the estimand framework would facilitate the implementation and advancement of the estimand framework in the pharmaceutical industry.